Tohban Report 2011-08-17
From RHESSI Wiki
|Start Date:||10 August 2011|
|End Date:||17 August 2011|
|Next Tohban:||Claire Raftery|
|List all reports|
After the excitement of last week's X class flare, activity this week diminished abruptly, with the base flux level dropping down to A class level. However, two AR (11271 and 11272) have recently rotated back onto the disk. These are productive regions that produced occulted flares even before their appearance on the disk and so we can expect that upcoming activity will be higher.
How many GOES flares occurred?
Flares above B, C, M, X class were 14 10 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 11 9 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 4 3 0 0
The SSR is fully emptying after each pass. Max fill was 39%.
There were no data gaps this week.
Detector 8 went haywire, producing long periods of very high noise in the count rates, starting on August 10. See this link for an example:
In an effort to mitigate this problem, the voltage to detector 8 was lowered by six steps (to 195) on Sunday (August
Hugh pointed out an anomaly in the spacecraft pointing, causing errors in flare locations. An example is here:
Martin and others are looking into this.